OpenStudy (anonymous):

what is the length of the arc on a circle with radius 20in intercepted by a 15 degree angle use pi=3.14

OpenStudy (mathmale):

What is the formula for "arc length" when the radius of a circle is given, as is the central angle that defines the arc?

OpenStudy (amity):

15/180 (pi r)=3.14(20)/12=62.8/12=5.23in